which of the quadratic functions has the widest graph?
a) y = 1/3 x^2
b) y = -4x^2
c) y = 0.3x^2
d) y = - 4/5x^2

Found 2 solutions by stanbon, advanced_Learner:
Answer by stanbon(75887)

which of the quadratic functions has the widest graph?
a) y = 1/3 x^2 = 0.333..x^2
b) y = -4x^2 = -4 x^2
c) y = 0.3x^2 = 0.3x^2
d) y = - 4/5x^2 = -0.8x^2
------
Widest means rises or descends at the slowest rate::
Ans:: y = 0.3x^2
------
Cheers,
Stan H.
---------------

Answer by advanced_Learner(501)

the one with the least coefficient has the widest area in this
y=0.3x^2